OpenStudy (ericanoel912):

FAN AND MEDAL! If f(x) = log2 (x + 4), what is f−1(3)?

OpenStudy (irishboy123):

\(y = \log_2 (x+4)\) \(2^y = x+4\) \(2^y -4 = x\)